OpenStudy (anonymous):

combine like terms and simplify for a medal and a fan! 1) 3n-13-5n+6n 2) 2rt+2r+7rt

OpenStudy (anonymous):

1) 4n-13 and

OpenStudy (anonymous):

erh.... 1) n(3-5+6)-13=4n -13 2) rt(2+7)+2r=r(t(2+7)+2)=r(9t+2)

OpenStudy (anonymous):

i think 2) 11r2t

OpenStudy (anonymous):

or 2) 9rt +2r

OpenStudy (anonymous):

I still do not understand

OpenStudy (anonymous):

@Syderitic

OpenStudy (anonymous):

ok, which part?

OpenStudy (anonymous):

@INeedHelpWithThisQuestion

OpenStudy (anonymous):

The answer

OpenStudy (anonymous):

@Syderitic

OpenStudy (anonymous):

did you understand 1)?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

didn't you just solve a related rates problem? -_-'? 2) anyway... i'm assuming rt is r times t and not unrelated to r. You begin by grouping like terms, terms afected by rt are 2rt and 7rt giving you 9rt +2r

OpenStudy (anonymous):

erh....you gave me a medal...so did you got it? @INeedHelpWithThisQuestion

OpenStudy (anonymous):

no i just did because you are being helpful

OpenStudy (anonymous):

but i got it now

OpenStudy (anonymous):

lol I'm so dumb i can't believe i did not know